# Benchmarking AI Models in Software Engineering: A Review, Search Tool, and Unified Approach for Elevating Benchmark Quality

### v2

@@ 3.2. Results of the Review
CRUXEval-X (Xu et al., 2025a) 19 PLs 19K
Merge Conflict Repair ConflictBench (Shen and Meng, 2024) Java 180
Type Inference TypeEvalPy  (Venkatesh et al., 2023) Python 845 (annotated labels)
TypeEvalPy AutoGen  (Venkatesh et al., 2023) Python 78373 (annotated labels)
Automatic Code Quality Review CodeReview  (Li et al., 2022b) 8 languages 7.9M pull requests
Software Maintainability  (Schnappinger et al., 2020) Java 519 projects
@@
BenMark (Wang et al., 2025d) Java 1,299,186 methods
CodeReviewer (Li et al., 2022c) 9 languages 13,100
CodeReview-New (Guo et al., 2024) 16 languages 14,600
CodeReviewQA (Lin et al., 2025) 9 languages 900
Hallucination Detection HALLUCODE (Liu et al., 2024) Python 5,663
CodeHaluEval (Tian et al., 2025) Python 699
Collu-Bench (Jiang et al., 2024a) N/A 13,234
LMDefects (Fan et al., 2023a) Java 60 (Easy), 53 (Medium)
Codemirage (Agarwal et al., 2024) Python 1,137
With AI4SE models mainly being utilized for program synthesis, it remains relatively questionable how effective these models are in generating tests and repairing bugs, as it is unclear whether these models truly understand code. For example, Siddiq et al. (Siddiq et al., 2024) observed Codex (Chen et al., 2021) being able to get above 80% coverage for HumanEval (Chen et al., 2021), yet many test smells were discovered and for another dataset, no higher than 2% coverage was attained. This reveals the importance of benchmarking AI4SE models’ capabilities in test generation, bug repair, and understanding. In Table 15, several benchmarks are listed that make an effort to assess the aforementioned. Additionally, Table 16 table features benchmarks that have been designed to evaluate a model’s capabilities in dealing with everyday tasks of a software engineer (e.g., merge-conflict repair, Code Reviews, etc.).
Table 17. Overview of Multi-Category Benchmarks, Covering Various Tasks.
@@
Big-Bench (Srivastava et al., 2022) Python, Numeric, JSON, English Functions over numbers, Mathematical Reasoning, Text2Code, Code2Text, Code explanation, Debugging, Turing Complete Concept Learning, amongst other tasks 250, several per category, 42, 60, 66, 34, 6,390
@@
XLCoST (Zhu et al., 2022a) C, C++, C#, Java, JavaScript, Kotlin, PHP, Python, Ruby, Rust Text2Code (program synthesis, code search), Code Summarization, Code Translation 567K (509k, 58k), 567K, 122K
@@
CrossCodeBench (Niu et al., 2023) Java, C#, Python, C++, JS, PHP, Go, Ruby, TS, C, Bash, Shell Classification, In-Filling, Translation, Generation, Summarization, Type Prediction, Question Answering 6.6M, 13.4M, 2.4M, 19.5M, 11.2M, 773K, 190K
@@
Long Code Arena (Bogomolov et al., 2024) English, Python, Java, Kotlin Commit Message Generation, Module Summarization, Library-Based Code Generation, Project-Level Code Completion, Bug Localization, CI Builds Repair 163, 216, 150, 908 (varying sizes), 14.96K, 78
@@
CodeXGLUE  (Bogomolov et al., 2024) MicrosoftDocs [^18] https://github.com/MicrosoftDocs/ CodeSearchNet (Husain et al., 2019) English, Chinese, Norwegian, Danish, Latvian Go, Java, JavaScript, PHP, Python, Ruby Code Documentation Translation, Code Documentation (Code Summarization, Comment Generation) (CN: 52K, NOR: 26K, DK: 45K, LT: 21K), 621870
@@
DomainEval (Zhu et al., 2025) Python Computation, Network, Basic operation, System, Visualization, Cryptography 5,892 cases total
@@
CodeApex (Liu et al., 2022) C++, English, Chinese Programming comprehension Code generation, Code correction 250, 476, 1330
@@
Additionally, wee collect a set of generic benchmarks in Table 17. These benchmarks are not only single-task benchmarks like others previously seen in this section but are rather a collection of tasks spanning a wide range of languages and task types. A notable, and widely known benchmark in this category is the BigBench Benchmark (Srivastava et al., 2022) which consists of 167 tasks (not all relevant to AI4SE).
@@
Table 18. Overview of Selected Benchmarks for Log Statement Generation and Parsing.
Category Benchmark Language(s) No. of Problems
Log Statement Generation LANCE (Mastropaolo et al., 2022) Java 76,421
LogBench (Li et al., 2023c) Java 6,849
SCLogger (Li et al., 2024b) Java 31,170
AL-Bench (Tan et al., 2025) Java 39,600
Log Parsing LogBase (Zhang et al., 2025d) Various Languages 85,300
LogHub (Zhu et al., 2023) Various Languages 32,000
LogHub-2.0 (Jiang et al., 2024b) Various Languages 3.6M
LogPM (Hashemi et al., 2024) Various Languages 10,821,589
LogEval (Cui et al., 2024) English, Chinese 4,000 log entries across 4 tasks
@@
Finally, we have identified a rather large count of papers that have to do with the matter of Log Parsing and Log Statement generation. we have presented these benchmarks in Table 18.
3.3. Limitations of Existing Benchmarks
Table 19. Common limitations in AI4SE benchmarks
Limitation Description Representative Examples
@@ 3.3. Limitations of Existing Benchmarks
Language Specific Focuses only on one programming language HumanEval, MBPP, APPS
@@
No Leaderboard No official leaderboard; hinders fair and consistent comparisons SPoC, RepoEval, CONCODE
@@
Poor Maintenance / Data Quality Dataset contains outdated problems, lacks documentation, or is not updated regularly ToolQA, ExampleCheck
@@
Not Peer Reviewed Not been published in a peer-reviewed venue CodeSearchNet, Methods2Test
@@
Infrequently Used Rarely cited in academic work JavaBench, CoIR
@@
Dataset Creation Method How problems were obtained: mined from real-world sources, created by humans, or LLM-generated AVATAR (mined), Spider (created), CRUXEval (generated)
@@
Table 19 summarizes these limitations and provides representative benchmarks for each. The full list of evaluated benchmarks and their metadata can be found in the replication package of the study.
In Conclusion, based on our in-depth inspection of HumanEval and MBPP and combined with the inspection of other benchmarks in our review, we obtained an overview of the limitations in current AI4SE benchmarks which we use as a guide in shaping our methodology proposed in Section 5.
4. BenchScout- Locating AI4SE Benchmarks
Due to the abundance of AI4SE benchmarks, identifying the most suitable one for a specific software engineering task can be challenging. As a result, many default to evaluating their models on popular benchmarks like HumanEval (Chen et al., 2021) which has its own flaws.
@@ 4. BenchScout- Locating AI4SE Benchmarks
4.1. Context Extraction and Visualization
4.1.1. Overview of the Semi-Automated Pipeline.
@@ 4.1. Context Extraction and Visualization
4.1.2. Balancing Automation and Oversight.
@@
4.1.3. Interactive Visualization.
The final layer of the pipeline provides an interactive 2D interface that enables users to explore benchmarks based on their semantic similarity. The visualization supports fuzzy search, local citation graphs, and interactive paper details. Importantly, user interactions on the frontend—such as suggesting new benchmarks—are routed back into the data collection layer (cf. Figure 5).
4.2. Additional Features
@@
Figure 5. Pipeline architecture of BenchScout. This includes automated metadata extraction, semantic embedding, clustering, and interactive visualization.
4.3. User Study
With all the mentioned features, we aim to create a platform that can be extended and used by both academics and practitioners alike to find the appropriate dataset/benchmark for their use case with more ease. To evaluate how effective and usable this new tool is for the end-users, we conducted a user study on 22 people from both industry (9) and academia (13). In selecting demographics for the BenchScout user study, we aimed to assess the tool’s effectiveness across a diverse group of users with varying degrees of expertise. This approach seeks to determine the tool’s applicability for individuals at either end of the spectrum—whether they are beginners or seasoned experts, from academia or industry. The ultimate objective of the tool is to facilitate the selection of the appropriate benchmark, making it more accessible irrespective of prior knowledge. For this, we had each participant interact with the tool for however long they saw fit and asked them to fill out a questionnaire consisting of eleven 5-point Likert Scale questions and three open questions.
4.3.1. Questionnaire Design
Table 20. Questionnaire Design Overview
@@ 4.3.1. Questionnaire Design
Participant Background What is your professional background?
What is your role?
How familiar are you with AI4SE benchmarks?
How many years of experience do you have in this field?
5-point Likert Scale
@@
Search Functionality How easy was it to navigate the interface?
How intuitive was the search functionality?
How effective was the tool in finding benchmarks?
Was the visual evaluation of datasets useful?
5-point Likert Scale
@@
User Experience & Feedback How would you rate the overall user experience?
How likely are you to use the tool in your work?
Did you experience any issues or challenges?
What other tools do you use for searching benchmarks?
How does this tool compare to others?
How well does the tool meet the needs of professionals in AI4SE?
Likert Scale, Open-ended Open-ended questions for in-depth feedback
The questionnaire designed for the study was divided into four key sections to gather feedback about the tool, namely the participant’s background, the quality of the search functionality, the quality of the cross-referencing feature, and the overall user experience. Table 20 is an overview of the posed questions.
4.3.2. Results and Analysis
@@
  3. (3)
    Familiarity with AI4SE: On a scale of 1 (Not familiar) to 5 (Very familiar), 8 participants rated themselves a ”5” and 7 rated themselves a ”4,” indicating a strongly informed participant pool
In terms of search functionality, the tool scored high on usability with an average rating of 4.50, showing that users found it easy to navigate. The intuitiveness of the search interface received a solid 4.1, while its effectiveness in helping users find benchmarks was rated 4.0. The visual evaluation feature received a rating of 3.8, indicating some room for improvement in how visual elements assist in the search process.
@@
In conclusion, while the tool is largely perceived as useful and user-friendly, there were several areas for improvement, particularly around visualization, citation handling, and filtering options, which could enhance the overall user experience and the tool’s effectiveness in AI4SE research. Given the time constraints, we prioritized and implemented key features, leaving some for the future.
5. BenchFrame
In this section, we propose BenchFrame, a detailed and peer-review-oriented framework for improving the quality of existing benchmarks. We explain our approach through a case study in which we propose HumanEvalNext, a corrected foundation based on the HumanEval benchmark.
The HumanEval benchmark has long been the de facto standard for evaluating the code-generation capabilities of AI4SE models. It has recently been used to evaluate the latest and greatest LLMs from large companies such as Google Gemini (Gemini Team, 2023) and OpenAI GPT4 (OpenAI, 2023). Despite the great fame, however, our review in Section 3 points towards the existence of numerous notable problems with this foundational benchmark, namely, the existence of incorrect tests, suboptimal canonical solutions, and imprecise problem definitions amongst many others.
5.1. Approach
Full Code Review of HumanEval Standardized Observations Check Observations HumanEval Variants
Modify HumanEval ( $\rightarrow$  HumanEvalNext  ) Peer-Review Changes Evaluate AI4SE Models
Figure 6. Outline of the general approach behind BenchFrame through a case-study of HumanEval.
To improve the quality of the given benchmark, we pursue the following approach – also illustrated in Figure 6. First, we initiate a comprehensive code review, leading to standardized observations (Section 5.1.1). Then, we address these identified issues through a series of modifications (Section 5.1.2), followed by a peer review (Section 5.1.3) to ensure accuracy and reliability. Finally, we experiment with the revised benchmark to evaluate and discuss the results. Below, more details are provided for specific steps in this approach.
5.1.1. Standardized Observations in Current HumanEval Benchmarks
@@ 5.1.2. Modifications in HumanEvalNext
In HumanEvalNext, we address all the above issues by manually modifying all problems in the original HumanEval benchmark. We decide to improve the original HumanEval as (1) flaws in the original version persist even in improved versions such as HumanEvalPlus, and (2) the fact that HumanEval is still widely used in new literature. While in our we have gone for the original benchmark, due to the adaptable nature of BenchFrame one could opt to have an improved version of the dataset as the starting point. To summarize, these are the general changes made in HumanEvalNext and their benefits: In this work, several key improvements have been made to address the shortcomings of previous benchmarks. First, all suboptimal and incorrect canonical solutions have been fixed, which were previously missed due to insufficient testing and a lack of comprehensive quality review. Furthermore, type annotations have been added to all problems, offering valuable context and simplifying the translation to other programming languages. The original HumanEval benchmark only included type annotations for the first 30 problems, representing merely 18% of the total set of 164 problems. In addition, better support for language conversion frameworks has been incorporated. For instance, HumanEvalNext now features improved compatibility with frameworks like MultiPL-E (Cassano et al., 2022), which supports translation to 18 additional programming languages by standardizing all tests to equality assertions where feasible. This change has reduced the number of incompatible problems by a factor of ten.
@@
Besides these adjustments, challenging scenarios (such as negative values, zero instances, empty inputs, and non-alphanumeric symbols) are incorporated into each task to guarantee that only top-quality AI4SE models, adept at addressing diverse situations, succeed. Accordingly, assertions are implemented within the code wherever constraints are detailed in the problem description. This measure prevents models from ignoring crucial details, thus improving the benchmark’s worth. In particular, we employ specification-based testing to assess functions; using boundary analysis, we explore combinations of within, on, and outside points. Furthermore, the test examples in the problem descriptions have been refined, which significantly affects model performance (Austin et al., 2021). Problems with excessive test examples now feature a reduced set, distributing the evaluation workload more fairly. Lastly, spelling errors have been corrected, descriptions have been consistently formatted, and problem descriptions have been aligned with the implementations, while still leaving room for models to demonstrate intuitive problem-solving skills expected of high-quality AI4SE models. The difficulty level has also been raised by incorporating more edge cases and modifying various problems, aiming to better reflect real-world challenges faced by engineers and to mitigate issues related to data leakage and saturated performance on the HumanEval leaderboards.
To illustrate the modifications of the tests in HumanEvalNext, consider Table 21.
@@
Table 21. Comparison of test statistics between HumanEval (based on https://github.com/openai/human-eval/blob/master/data/HumanEval.jsonl.gzhuman-eval-v2-20210705.json) and HumanEvalNext.
Metric HumanEval (Original) HumanEvalNext $\Delta$
@@
Average number of asserts 8 16 $\times 2$
Median number of asserts 7 11 $\times 1.57$
@@
Problems with ¡ 5 asserts 34 2 -94%
5.1.3. Peer Review Process of HumanEvalNext

### v3

@@ III-B Results of the Review
CRUXEval-X [218] 19 PLs 19K
Merge Conflict Repair ConflictBench [219] Java 180
Type Inference TypeEvalPy  [220] Python 845 (annotated labels)
TypeEvalPy AutoGen  [220] Python 78373 (annotated labels)
Automatic Code Quality Review CodeReview  [221] 8 languages 7.9M pull requests
Software Maintainability  [222] Java 519 projects
@@
BenMark [223] Java 1,299,186 methods
CodeReviewer [224] 9 languages 13,100
CodeReview-New [225] 16 languages 14,600
CodeReviewQA [226] 9 languages 900
Hallucination Detection HALLUCODE [227] Python 5,663
CodeHaluEval [228] Python 699
Collu-Bench [229] N/A 13,234
LMDefects [230] Java 60 (Easy), 53 (Medium)
Codemirage [231] Python 1,137
With AI4SE models mainly being utilized for program synthesis, it remains relatively questionable how effective these models are in generating tests and repairing bugs, as it is unclear whether these models truly understand code. For example, Siddiq et al. [232] observed Codex [1] being able to get above 80% coverage for HumanEval [1], yet many test smells were discovered and for another dataset, no higher than 2% coverage was attained. This reveals the importance of benchmarking AI4SE models’ capabilities in test generation, bug repair, and understanding. In Table XV, several benchmarks are listed that make an effort to assess the aforementioned. Additionally, Table XVI table features benchmarks that have been designed to evaluate a model’s capabilities in dealing with everyday tasks of a software engineer (e.g., merge-conflict repair, Code Reviews, etc.).
TABLE XVII: Overview of Multi-Category Benchmarks, Covering Various Tasks.
@@
Big-Bench [233] Python, Numeric, JSON, English Functions over numbers, Mathematical Reasoning, Text2Code, Code2Text, Code explanation, Debugging, Turing Complete Concept Learning, amongst other tasks 250, several per category, 42, 60, 66, 34, 6,390
@@
XLCoST [234] C, C++, C#, Java, JS, Kotlin, PHP, Python, Ruby, Rust Text2Code (program synthesis, code search), Code Summarization, Code Translation 567K (509k, 58k), 567K, 122K
@@
CrossCodeBench [235] Java, C#, Python, C++, JS, PHP, Go, Ruby, TS, C, Bash, Shell Classification, In-Filling, Translation, Generation, Summarization, Type Prediction, Question Answering 6.6M, 13.4M, 2.4M, 19.5M, 11.2M, 773K, 190K
@@
Long Code Arena [236] English, Python, Java, Kotlin Commit Message Generation, Module Summarization, Library-Based Code Generation, Project-Level Code Completion, Bug Localization, CI Builds Repair 163, 216, 150, 908 (varying sizes), 14.96K, 78
@@
CodeXGLUE  [236] MicrosoftDocs [^18] https://github.com/MicrosoftDocs/ CodeSearchNet [237] English, Chinese, Norwegian, Danish, Latvian Go, Java, JS, PHP, Python, Ruby Code Documentation Translation, Code Documentation (Code Summarization, Comment Generation) (CN: 52K, NOR: 26K, DK: 45K, LT: 21K), 621870
@@
DomainEval [238] Python Computation, Network, Basic operation, System, Visualization, Cryptography 5,892 cases total
@@
CodeApex [23] C++, English, Chinese Programming comprehension Code generation, Code correction 250, 476, 1330
@@
Additionally, we collect a set of generic benchmarks in Table XVII. These benchmarks are not only single-task benchmarks like others previously seen in this section but are rather a collection of tasks spanning a wide range of languages and task types. A notable, and widely known benchmark in this category is the BigBench Benchmark [233] which consists of 167 tasks (not all relevant to AI4SE).
@@
TABLE XVIII: Overview of Selected Benchmarks for Log Statement Generation and Parsing.
Category Benchmark Language(s) #Problems
Log Statement Generation LANCE [239] Java 76,421
LogBench [240] Java 6,849
SCLogger [241] Java 31,170
AL-Bench [242] Java 39,600
Log Parsing LogBase [243] Various 85,300
LogHub [244] Various 32,000
LogHub-2.0 [245] Various 3.6M
LogPM [246] Various 10,821,589
LogEval [247] Eng., Chin. 4K logs (4 tasks)
@@
Finally, we have identified a rather large count of papers that have to do with the matter of Log Parsing and Log Statement generation. we have presented these benchmarks in Table XVIII.
III-C Limitations of Existing Benchmarks
TABLE XIX: Common limitations in AI4SE benchmarks
Limitation Description Representative Examples Frequency Observed
@@ III-C Limitations of Existing Benchmarks
Language Specific Focuses only on one programming language HumanEval, MBPP, APPS 169
@@
No Leaderboard No official leaderboard; hinders fair and consistent comparisons SPoC, RepoEval, CONCODE 201
@@
Poor Maintenance/Data Quality Dataset contains outdated problems, lacks documentation, or is not updated regularly ToolQA, ExampleCheck 218
@@
Not Peer Reviewed Not been published in a peer-reviewed venue CodeSearchNet, Methods2Test 79
@@
Infrequently Used Rarely cited in academic work JavaBench, CoIR 199
@@
Dataset Creation Method How problems were obtained: mined from real-world sources, created by humans, or LLM-generated AVATAR (mined), Spider (created), CRUXEval (generated) Mined: 149, Created: 69, Generated: 46
@@
Table XIX summarizes these limitations and provides representative benchmarks for each. The full list of evaluated benchmarks and their metadata can be found in the replication package of the study.
In conclusion, based on our in-depth inspection of HumanEval and MBPP and combined with the inspection of other benchmarks in our review, we obtained an overview of the limitations in current AI4SE benchmarks which we use as a guide in shaping our methodology proposed in section V.
IV BenchScout- Locating AI4SE Benchmarks
Due to the abundance of AI4SE benchmarks, identifying the most suitable one for a specific SE task can be challenging. As a result, many default to evaluating their models on popular benchmarks like HumanEval [1] which has its own flaws.
@@ IV BenchScout- Locating AI4SE Benchmarks
IV-A Context Extraction and Visualization
IV-A1 Overview of the Semi-Automated Pipeline.
@@ IV-A Context Extraction and Visualization
IV-A2 Balancing Automation and Oversight.
@@
IV-A3 Interactive Visualization.
The final layer of the pipeline provides an interactive 2D interface that enables users to explore benchmarks based on their semantic similarity. The visualization supports fuzzy search, local citation graphs, and interactive paper details. Importantly, user interactions on the frontend, such as suggesting new benchmarks, are routed back into the data collection layer (cf. Figure 5).
IV-B Additional Features
@@
Figure 5: Pipeline architecture of BenchScout. This includes automated metadata extraction, semantic embedding, clustering, and interactive visualization.
IV-C User Study
With all the mentioned features, we aim to create a platform that can be extended and used by both academics and practitioners alike to find the appropriate dataset/benchmark for their use case with more ease. To evaluate how effective and usable this new tool is for the end-users, we conducted a user study on 22 people from both industry (9) and academia (13). In selecting demographics for the BenchScout user study, we aimed to assess the tool’s effectiveness across a diverse group of users with varying degrees of expertise. This approach seeks to determine the tool’s applicability for individuals at either end of the spectrum; whether they are beginners or seasoned experts, from academia or industry. The ultimate objective of the tool is to facilitate the selection of the appropriate benchmark, making it more accessible irrespective of prior knowledge. For this, we had each participant interact with the tool for however long they saw fit and asked them to fill out a questionnaire consisting of eleven 5-point Likert Scale questions and three open questions.
IV-C1 Questionnaire Design
TABLE $X$: Questionnaire Design Overview
@@ IV-C1 Questionnaire Design
Participant Background What is your professional background? What is your role? How familiar are you with AI4SE benchmarks? How many years of experience do you have in this field? 5-point Likert Scale
@@
Search Functionality How easy was it to navigate the interface? How intuitive was the search functionality? How effective was the tool in finding benchmarks? Was the visual evaluation of datasets useful? 5-point Likert Scale
@@
User Experience & Feedback How would you rate the overall user experience? How likely are you to use the tool in your work? Did you experience any issues or challenges? What other tools do you use for searching benchmarks? How does this tool compare to others? How well does the tool meet the needs of professionals in AI4SE? Likert Scale, Open-ended Open-ended questions for in-depth feedback
The questionnaire designed for the study was divided into four key sections to gather feedback about the tool, namely the participant’s background, the quality of the search functionality, the quality of the cross-referencing feature, and the overall user experience. Table XX is an overview of the posed questions.
IV-C2 Results and Analysis
@@
  3. 3.
    Familiarity with AI4SE: On a scale of 1 (Not familiar) to 5 (Very familiar), 8 participants rated themselves a “5” and 7 rated themselves a “4,” indicating a strongly informed participant pool
In terms of search functionality, the tool scored high on usability with an average rating of 4.5, showing that users found it easy to navigate. The intuitiveness of the search interface received a solid 4.1, while its effectiveness in helping users find benchmarks was rated 4.0. The visual evaluation feature received 3.8, indicating some room for improvement in how visual elements assist in the search process.
@@
In conclusion, the tool is largely perceived as useful and user-friendly, though several areas, particularly around visualization, citation handling, and filtering options, could be improved to enhance the overall user experience and the tool’s effectiveness. Due to time constraints, we prioritized and implemented key features, leaving some for the future.
V BenchFrame
In this section, we propose BenchFrame, a detailed and peer-review-oriented approach for improving the quality of existing benchmarks. We explain our approach through a case study in which we propose HumanEvalNext, a corrected foundation based on the HumanEval benchmark.
The HumanEval benchmark has long been the de facto standard for evaluating the code-generation capabilities of AI4SE models. It has recently been used to evaluate the latest and greatest LLMs from large companies such as Google Gemini [2] and OpenAI GPT4 [3]. Despite the great fame, however, our review in section III points towards the existence of numerous notable problems with this foundational benchmark, namely, the existence of incorrect tests, suboptimal canonical solutions, and imprecise problem definitions amongst others.
V-A Approach
Full Code Review of HumanEval Standardized Observations Check Observations HE Variants Modify HumanEval ($\rightarrow$ HumanEvalNext) Peer-Review Changes Evaluate AI4SE Models
Figure 6: BenchFrame’s approach through a case-study of HumanEval.
To improve the quality of the given benchmark, we pursue the following approach, also illustrated in Figure 6. First, we initiate a comprehensive code review, leading to standardized observations (subsubsection V-A1). Then, we address these identified issues through a series of modifications (subsubsection V-A2), followed by a peer review (subsubsection V-A3) to ensure accuracy and reliability. Finally, we experiment with the revised benchmark to evaluate and discuss the results. Below, more details are provided for specific steps in this approach.
V-A1 Standardized Observations in Current HumanEval Benchmarks
@@ V-A2 Modifications in HumanEvalNext
In HumanEvalNext, we address all the above issues by manually modifying all problems in the original HumanEval benchmark. We decide to improve the original HumanEval as (1) flaws in the original version persist even in improved versions such as HumanEvalPlus, and (2) the fact that HumanEval is still widely used in new literature. While in our we have gone for the original benchmark, due to the adaptable nature of BenchFrame one could opt to have an improved version of the dataset as the starting point. To summarize, these are the general changes made in HumanEvalNext and their benefits: In this work, several key improvements have been made to address the shortcomings of previous benchmarks. First, all suboptimal and incorrect canonical solutions have been fixed, which were previously missed due to insufficient testing and a lack of comprehensive quality review. Furthermore, type annotations have been added to all problems, offering valuable context and simplifying the translation to other programming languages. The original HumanEval benchmark only included type annotations for the first 30 problems, representing merely 18% of the total set of 164 problems. In addition, better support for language conversion frameworks has been incorporated. For instance, HumanEvalNext now features improved compatibility with frameworks like MultiPL-E [6], which supports translation to 18 additional programming languages by standardizing all tests to equality assertions where feasible. This change has reduced the number of incompatible problems by a factor of ten.
@@
Besides these adjustments, challenging scenarios (such as negative values, zero instances, empty inputs, and non-alphanumeric symbols) are incorporated into each task to guarantee that only top-quality AI4SE models, adept at addressing diverse situations, succeed. Accordingly, assertions are implemented within the code wherever constraints are detailed in the problem description. This measure prevents models from ignoring crucial details, thus improving the benchmark’s worth. In particular, we employ specification-based testing to assess functions; using boundary analysis, we explore combinations of within, on, and outside points. Furthermore, the test examples in the problem descriptions have been refined, which significantly affects model performance [15]. Problems with excessive test examples now feature a reduced set, distributing the evaluation workload more fairly. Lastly, spelling errors have been corrected, descriptions have been consistently formatted, and problem descriptions have been aligned with the implementations, while still leaving room for models to demonstrate intuitive problem-solving skills expected of high-quality AI4SE models. The difficulty level has also been raised by incorporating more edge cases and modifying various problems, aiming to better reflect real-world challenges faced by engineers and to mitigate issues related to data leakage and saturated performance on the HumanEval leaderboards. To illustrate the modifications of the tests in HumanEvalNext, consider Table XXI.
@@
TABLE XXI: Comparison of test statistics between HumanEval (based on https://github.com/openai/human-eval/blob/master/data/HumanEval.jsonl.gzhuman-eval-v2-20210705.json) and HumanEvalNext.
Metric HumanEval HumanEvalNext $\Delta$
@@
Avg. number of asserts 8 16 $\times 2$
Med. number of asserts 7 11 $\times 1.57$
@@
Total with << 5 asserts 34 2 -94%
V-A3 Peer Review Process of HumanEvalNext